V 9. Solve: y +y = t sint, y(0) = 0, y(0) = 1.

Hello
could you please solve this problem with the clear hands writing to read it please ? Also the good explanation to understand the solution is by step by step please

the subject is Laplace Transforms

the course is Complex analysis

thank you
Answer #1

9. y+y = Hisint, y0 = 9 70 = 1 Auxilarry equation mrt1=0 mati complementary fum y (t) a cost the sint particular integral Ieit 21 2Di eit 21 teit sint = (-1) (esat +181nt) (-1) imaginary part of 3 [Feat Literat titsint-& tsint 2 21 [ + cost-it sintgiven Condelfons y@=0 4 = 0 and ya) = 1 0 +62 = 1 so y t = sint- t 4 ť lost & t sint + ( - 2 + 1) sint cost & Anos)For any queries comment...